0.514g of gas occupies 150.0ml at 23.0 degrees celsius and 755 torr. What is its molar mass? What gas is it?
Answer #1

m = 0.514

V = 150 ml or .15 L

T = 23 or 296 K

p = 755 torr or 0.993 atm

find MW

first find moles

PV = nRT

n = PV/(RT) = (0.993)(0.15)/(0.082*296) = 0.006136 gmol

find MW

MW = mass/mol = 0.514 / 0.006136 = 83.75 g/gmol

could be Hydrogen Bormide which MW = 81 g/gmol